OpenStudy (anonymous):

rewrite√8 in index form?

OpenStudy (anonymous):

\[\sqrt{2^3}\] \[=2^{\frac{3}{2}}\]

OpenStudy (anonymous):

my text gives the answer 8 1/2?????

OpenStudy (anonymous):

sqrt(x)=x^1/2

OpenStudy (anonymous):

so sqrt(8)=8^1/2

OpenStudy (anonymous):

I didn't know what index form was :-P

OpenStudy (anonymous):

i do not get that

OpenStudy (anonymous):

What do you mean?

OpenStudy (turingtest):

index form is an unusual terminology. it just shows what that radical signs are really fractional exponents, so the 5th root of X is X^1/5, or the 3rd root of Y^4 is Y^(4/3) etc.

OpenStudy (turingtest):

make sense?

OpenStudy (anonymous):

sort of

OpenStudy (anonymous):

i get it